# Decomposition of the Tschirnhausen module for coverings on decomposable $\mathbb{P}^{1}$-bundles

### v1

@@ Abstract.
In this note, we show that for a smooth algebraic variety $Y$ and a smooth $m$-secant section $X$ of the $\mathbb{P}^{1}$-bundle
@@ 1. Introduction and statement
is a smooth, irreducible divisor (an $m$-secant section of $f$), and let $\varphi:=f|_{X}:X\longrightarrow Y$ be the restriction of the projection $f$ to $X$. Then:
@@
    $\varphi_{\ast}\mathcal{O}_{X}\!\big{(}(H-f^{\ast}E)|_{X}\big{)}\;\cong\;\mathcal{O}_{Y}\;\oplus\;\mathcal{O}_{Y}(-E)\;\oplus\;\mathcal{O}_{Y}(-E-\Delta)\;\oplus\;\cdots\;\oplus\;\mathcal{O}_{Y}\bigl{(}-(m-1)E-\Delta\bigr{)}.$
@@ 2. Proof of Theorem 1
$0\to\mathcal{O}_{S}\bigl{(}-(m-1)H-f^{\ast}(E+\Delta)\bigr{)}\to\mathcal{O}_{S}(H-f^{\ast}E)\to\mathcal{O}_{X}\bigl{(}(H-f^{\ast}E)_{|_{X}}\bigr{)}\to 0,$
@@
    $f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H-f^{\ast}(E+\Delta)\bigr{)}\cong f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H\bigr{)}\otimes\mathcal{O}_{Y}(-E-\Delta)=0$,
@@
    $f_{\ast}\mathcal{O}_{S}(H-f^{\ast}E)\cong f_{\ast}\mathcal{O}_{S}(H)\otimes\mathcal{O}_{Y}(-E)\cong\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}$, and
@@
    $R^{1}f_{\ast}\mathcal{O}_{S}(H-f^{\ast}E)\cong R^{1}f_{\ast}\mathcal{O}_{S}(1)\otimes\mathcal{O}_{Y}(-E)=0$.
@@
(2) $0\to\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\to\varphi_{\ast}\mathcal{O}_{X}\bigl{(}(H-f^{\ast}E)_{|_{X}}\bigr{)}\to R^{1}f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H-f^{\ast}(E+\Delta)\bigr{)}\to 0.$
For the term $R^{1}f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H-f^{\ast}(E+\Delta)\bigr{)}$, we find:
$R^{1}f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H-f^{\ast}(E+\Delta)\bigr{)}$ $\cong R^{1}f_{\ast}\mathcal{O}_{S}\bigl{(}-(m-1)H\bigr{)}\otimes\mathcal{O}_{Y}(-(E+\Delta))$
@@
Computing $\operatorname{Ext}^{1}\bigl{(}R^{1}f_{\ast}\mathcal{O}_{S}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr{)}$, we obtain
$\operatorname{Ext}^{1}$ $\bigl{(}R^{1}f_{\ast}\mathcal{O}_{S}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr{)}$
@@
$\cong H^{1}\bigl{(}Y,\left[\mathcal{O}_{Y}(E)\oplus\oplus^{2}_{1}\mathcal{O}_{Y}(E)\oplus\cdots\oplus\oplus^{2}_{1}\mathcal{O}_{Y}((m-1)E)\oplus\mathcal{O}_{Y}(mE)\right]\otimes\mathcal{O}_{Y}(\Delta)\bigr{)}=0,$
@@ 3. The case of curves
In this section we specialize to the case where $Y$ is a smooth projective curve and $E$ is an effective nonspecial divisor of degree $e$ on $Y$. In this situation, the bundle
@@
The surface $\mathcal{B}$ contains a section $\sigma_{0}$ corresponding to the exact sequence
$\sigma_{0}:\quad 0\longrightarrow\mathcal{O}_{Y}\longrightarrow\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\longrightarrow\mathcal{O}_{Y}(-E)\longrightarrow 0.$
Let $Y_{0}:=\sigma_{0}(Y)$. Then $Y_{0}$ is the section of minimal self-intersection, satisfying $Y_{0}^{2}=-e$. The surface $\mathcal{B}$ also contains another section $\sigma_{1}$, associated with the exact sequence
$\sigma_{1}:\quad 0\longrightarrow\mathcal{O}_{Y}\longrightarrow\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\longrightarrow\mathcal{O}_{Y}(E)\longrightarrow 0,$
which corresponds to the surjection $\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\twoheadrightarrow\mathcal{O}_{Y}$. If we denote $Y_{1}:=\sigma_{1}(Y)$, then
@@
where $H$ is the tautological divisor as in the general case. For the intersection numbers we obtain
@@
    cut from $F$ by a degree $m$ hyperplane in $\mathbb{P}^{R}$;
@@
    If $P\notin X_{m}$, then $\phi$ is an m:1m\!:\!1 cover, and
@@
    If $P\in X_{m}$ and the tangent line to $X_{m}$ at $P$ meets $X_{m}$ with multiplicity two, then $\phi$ is an (m1):1(m-1)\!:\!1 cover, and:
@@
for some line bundles $L$ and $M$ on $Y$. It would be interesting to find examples of such decompositions where $X$ is not an $m$-secant section of a decomposable $\mathbb{P}^{1}$-bundle over $Y$.

### v3

@@ Abstract.
In this note, we show that for a smooth algebraic variety $Y$ and a smooth $m$-section $X$ of the $\mathbb{P}^{1}$-bundle
@@ 1. Introduction and statement
is a smooth, irreducible divisor (an $m$-section of $f$), and let $\varphi:=f|_{X}:X\longrightarrow Y$ be the restriction of the projection $f$ to $X$. Then:
@@
    $\varphi_{\ast}\mathcal{O}_{X}\!\big((H-f^{\ast}E)|_{X}\big)\;\cong\;\mathcal{O}_{Y}\;\oplus\;\mathcal{O}_{Y}(-E)\;\oplus\;\mathcal{O}_{Y}(-2E-\Delta)\;\oplus\;\cdots\;\oplus\;\mathcal{O}_{Y}\bigl(-(m-1)E-\Delta\bigr).$
@@ 2. Proof of Theorem 1
$0\to\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\to\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\to\mathcal{O}_{X}\bigl((H-f^{\ast}E)_{|_{X}}\bigr)\to 0,$
@@
    $f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\cong f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H\bigr)\otimes\mathcal{O}_{Y}(-E-\Delta)=0$,
@@
    $f_{\ast}\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\cong f_{\ast}\mathcal{O}_{\mathcal{B}}(H)\otimes\mathcal{O}_{Y}(-E)\cong\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}$, and
@@
    $R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\cong R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(1)\otimes\mathcal{O}_{Y}(-E)=0$.
@@
(2) $0\to\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\to\varphi_{\ast}\mathcal{O}_{X}\bigl((H-f^{\ast}E)_{|_{X}}\bigr)\to R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\to 0.$
For the term $R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)$, we find:
$R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)$ $\cong R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H\bigr)\otimes\mathcal{O}_{Y}(-(E+\Delta))$
@@
Computing $\operatorname{Ext}^{1}\bigl(R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr)$, we obtain
$\operatorname{Ext}^{1}$ $\bigl(R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr)$
@@
$\cong H^{1}\bigl(Y,\left[\mathcal{O}_{Y}(E)\oplus\oplus^{2}_{1}\mathcal{O}_{Y}(2E)\oplus\cdots\oplus\oplus^{2}_{1}\mathcal{O}_{Y}((m-2)E)\oplus\mathcal{O}_{Y}\bigl((m-1)E\bigr)\right]\otimes\mathcal{O}_{Y}(\Delta)\bigr)$
$=0,$
@@ 3. The case of curves
In this section, we specialize to the case where $Y$ is a smooth projective curve and $E$ is an effective, nonspecial divisor of degree $e$ on $Y$. In this situation, the bundle
@@
The surface $\mathcal{B}$ contains a section $\sigma_{0}$ corresponding to
$\sigma_{0}:\quad\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\twoheadrightarrow\mathcal{O}_{Y}.$
Let $Y_{0}:=\sigma_{0}(Y)$. Then $Y_{0}$ is the section of minimal self-intersection, satisfying $Y_{0}^{2}=-e$. The surface $\mathcal{B}$ also contains another section $\sigma_{1}$, associated with the surjection
$\sigma_{1}:\quad\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\twoheadrightarrow\mathcal{O}_{Y}(E).$
If we denote $Y_{1}:=\sigma_{1}(Y)$, then
@@
where $H$ is the tautological divisor as in the general case. The intersection numbers are given by
@@
    cut from $F$ by a degree $m$ hypersurface in $\mathbb{P}^{R}$;
@@
    If $P\notin X_{m}$, then $\phi$ is an m:1m\!:\!1 covering, and
@@
    If $P\in X_{m}$, $m\geq 3$, and the tangent line to $X_{m}$ at $P$ meets $X_{m}$ with multiplicity two, then $\phi$ is an (m1):1(m-1)\!:\!1 covering, and:
@@
for some line bundles $L$ and $M$ on $Y$. It would be interesting to find examples of such decompositions where $X$ is not an $m$-section of a decomposable $\mathbb{P}^{1}$-bundle over $Y$.